OpenStudy (howard-wolowitz):

If an object is dropped from a height, its downward speed theoretically increases linearly over time because the object is subject to the steady pull of gravity. Here are observational data on the speed of a ball dropped from a certain height at time x = 0: Time (seconds) X 0 0.2 0.4 0.6 0.8 Speed (m/sec) Y 0 1.92 3.58 6.01 7.88

OpenStudy (studygurl14):

Sorry, don't know

OpenStudy (howard-wolowitz):

Ok thank you!

OpenStudy (michele_laino):

it is simple, we have to compute this value: \[\Large r = 1.92 - \left( { - 0.992 + 9.925 \cdot 0.2} \right) = ...?\]

OpenStudy (howard-wolowitz):

ok one second plz

OpenStudy (michele_laino):

oops.. I have made a typo, here is the right formula: \[\Large r = 1.92 - \left( { - 0.092 + 9.925 \cdot 0.2} \right) = ...?\]

OpenStudy (howard-wolowitz):

i got 0.027

OpenStudy (michele_laino):

that's right!! :)